# Efficient Model-Stealing Attacks Against Inductive Graph Neural Networks

### v2

@@ Abstract
Graph Neural Networks (GNNs) are recognized as potent tools for processing real-world data organized in graph structures. Especially inductive GNNs, which enable the processing of graph-structured data without relying on predefined graph structures, are gaining importance in an increasingly wide variety of applications. As these networks demonstrate proficiency across a range of tasks, they become lucrative targets for model-stealing attacks where an adversary seeks to replicate the functionality of the targeted network. Large efforts have been made to develop model-stealing attacks that extract models trained with images and texts. However, little attention has been paid to stealing GNNs trained on graph data. This paper introduces a novel method for unsupervised model-stealing attacks against inductive GNNs, based on graph contrasting learning and spectral graph augmentations to efficiently extract information from the targeted model. The proposed attack is thoroughly evaluated on six datasets and the results show that our approach outperforms the current state-of-the-art by Shen et al. (2021). More specifically, our attack surpasses the baseline across all benchmarks, attaining superior fidelity and downstream accuracy of the stolen model while necessitating fewer queries directed toward the target model.
@@ 1 Introduction
Graph Neural Networks (GNNs) are designed to operate on graph-structured data, such as molecules or social networks [7, 37]. Unlike models designed for images and text, GNNs work by simultaneously considering both node features and graph structures. GNNs gain significant attention due to their effectiveness in tasks that are used as input structured data, including social networks. The advantage of GNNs is that they are designed to be scalable to large graphs while ensuring computational efficiency. On the other hand, it is also known that Machine Learning (ML) models are susceptible to model stealing attacks [31, 13, 5], where an adversary, with query access to a target (victim) model, can extract its parameters or functionality, often at a fraction of the cost for training from scratch. Therefore, the adversary first crafts a number of queries as the input to the target model’s API and obtains the corresponding outputs. Then, a local surrogate model is trained to match the output of the target for the same input data. As a consequence, the surrogate model may both violate the intellectual property of the target model and serve as a stepping stone for further attacks like extracting private training data [21] and constructing adversarial examples [25].
@@
Figure 1: Existing model stealing methods rely on aligning only the predictions for individual graph nodes sent to the victim and surrogate model. Our method greatly increases the amount of information extracted from the victim model outputs. We align the node representation from the surrogate model with the corresponding representation from the victim model while simultaneously distinguishing it from the representation of other nodes. By integrating spectral augmentations to augment the graph inputs of the surrogate models during its training process, we further increase the efficiency of the stealing process per query sent to the victim model.
@@
It is worth stressing that the majority of the current efforts on model stealing attacks concentrate on ML models for images and text data [31, 23, 34, 17]. Despite the growing value and importance of GNNs, model-stealing attacks against these types of neural networks are severely understudied, when compared to image or text models. To our knowledge, there exists only a single work [30] that targets inductive GNNs, i.e. a type of model that can generalize well to unseen nodes. The authors define the threat model for model stealing attacks on inductive GNNs and describe six attack methods based on the adversary’s background knowledge and the responses of the target models. During the attack, the response of the surrogate model is aligned with the response of the target model by minimizing the RMSE loss between them. Although their approach yields promising results, the underlying method of training the surrogate model does not efficiently utilize the information available from the target’s responses. This increases the number of queries required to extract the target model’s API which results in higher chances of attack detection and mitigation.
@@
Inspired by recent approaches to encoder model stealing [5, 20, 28] we leverage the significant fact that augmented graph inputs should produce similar model outputs. Thus, the proposed GNN model-stealing attack utilizes a contrastive objective at the node level. Specifically, for an input graph, we generate two graph views—one from the target model and one from the surrogate model for an augmented graph input. The surrogate model is trained by maximizing the agreement of node representations in these two views. This contrastive objective considers both positive and negative pairs between the node representations from the target and surrogate model. To augment the graph input to the surrogate model we leverage graph transformation operations derived via graph spectral analysis, enabling contrastive learning to capture useful structural representations. Those transformations include spectral graph cropping and graph frequency components reordering. Thus, the proposed methods that combine contrastive learning loss and effective graph-specific augmentations enable efficient usage of the information from the target model outputs. As a result, the proposed approach fits into the framework established by Shen et al. (2021) [30]. It presents an enhanced surrogate model training methodology that significantly augments the efficiency and performance of the model-stealing process, outperforming the state-of-the-art method with less than half of its query budget.
@@
In this paper, we evaluate all proposed attacks on three popular inductive GNN models including GraphSAGE [10], Graph Attention Network (GAT) [33], and Graph Isomorphism Network (GIN) [36] with six benchmark datasets, i.e., DBLP [24], Pubmed [27], Citeseer Full [8], Coauthor Physics [29], ACM [35], and Amazon Co-purchase Network for Photos [22]. The conducted extensive experiments demonstrate that our model-stealing attack consistently outperforms the state-of-the-art baseline approach on an overwhelming majority of datasets and models. It turned out that in the particular case of using SAGE as the surrogate model the increase in accuracy reached 36.8 %pt. for a GIN target trained on Amazon Co-purchase Network for Photos.
@@
  • We propose a new unsupervised GNN model stealing attack trained with a contrastive objective at the node level that compels the embedding of corresponding nodes from the surrogate and victim models to align while also distinguishing them from embeddings of other nodes within the surrogate and the target model.
  • Spectral graph extensions have been adapted to generate multiple graph views for the surrogate model during the contrastive learning process.
  • The broad comparison of the new stealing technique to the state of the art. The proposed stealing process shows better performance while requiring fewer queries against the target model.
@@ 2 Preliminaries
$u\in\mathcal{N}(v)$. Note that the node embedding at the $k$-th layer captures the information within its $k$-hop neighborhood.
@@
The notation applied in this paper is summarizåed in Table 1.
@@ 3 Threat Model
This section is dedicated to the description of the threat model, outlining the attack setting and the adversary’s goal and capabilities. The proposed approach is based on the threat model widely presented in [30].
Attack Setting. We operate within a challenging black-box setting in which the adversary lacks knowledge about the target GNN model, including model parameters and architecture, and cannot manipulate its training process, (e.g., training graph $\mathbf{G}_{O}$). Our paper specifically focuses on three node-level query responses, with the target model being an inductive GNN that takes node $v$’s $l$-hop subgraph $\mathbf{G}_{v}^{l}$ as input and provides the corresponding response for node $v$. This response could be the predicted posterior probability, node embedding vector, or a 2-dimensional t-SNE projection.
Adversary’s Goal. Following the taxonomy of [13], adversaries can pursue two goals: theft or reconnaissance. The theft adversary aims to construct a surrogate model $\mathcal{M}_{S}$ matching $\mathcal{M}_{T}$ on the target task [31, 25], jeopardizing intellectual property and violating confidentiality. In contrast, the reconnaissance adversary seeks a surrogate model $\mathcal{M}_{S}$ mirroring $\mathcal{M}_{T}$ across all inputs. This high-fidelity match serves as a strategic tool for subsequent attacks, such as crafting adversarial examples without direct queries to $\mathcal{M}_{T}$ [25].
Adversary’s Capabilities. Firstly, we assume that the adversary queries a target model $\mathcal{M}_{T}$ hidden behind a publicly accessible API [31, 23, 11, 12] and obtains responses $\mathbf{R}$ based on an input query graph $\mathbf{G}_{Q}$. These responses can be in the form of a node embedding matrix ($\mathbf{H}$), a predicted posterior probability matrix ($\boldsymbol{\Theta}$), or a t-SNE projection matrix of $\mathbf{H}$ ($\boldsymbol{\Upsilon}$), which represents common API outputs in real-world scenarios. This diversity in responses reflects different levels of knowledge an adversary might gain access to, spanning applications such as graph visualization, transfer learning, and fine-tuning pre-trained GNNs.
Secondly, we assume that the query graph $\mathbf{G}_{Q}$ (including node features $\mathbf{X}_{Q}$ and graph structure $\mathbf{A}_{Q}$) is drawn from the same distribution as the graph $\mathbf{G}_{O}$ used for training $\mathcal{M}_{T}$. In practice, we consider that $\mathbf{G}_{Q}$ and $\mathbf{G}_{O}$ come from the same distribution if they are randomly sampled from the same dataset. This assumption aligns with recent attacks on neural networks using parts of public datasets [13, 11], especially in graph-rich domains like social networks and molecular graphs. This assumption can be relaxed, as described in Section 4.2.
@@ 4 Attack Framework
This section aims to discuss attack scenarios that can be launched by the adversary depending on different levels of knowledge and describe the general framework of GNN model-stealing. This work adopted both, the attack structure and taxonomy proposed by Shen and co-authors [30].
@@ 4.1 Attack Taxonomy
The adversary relies on two key pieces of information for launching the model stealing attack: the query graph $\mathbf{G}_{Q}=(\mathbf{A}_{Q},\mathbf{X}_{Q},\mathbf{C}_{Q})$ and its corresponding response $\mathbf{R}$ (node embedding matrix $\mathbf{H}$, predicted posterior probability matrix $\boldsymbol{\Theta}$, or t-SNE projection matrix $\boldsymbol{\Upsilon}$). It’s important to note that the adversary might lack structural information for the query graph $\mathbf{G}_{Q}$ (e.g., missing adjacency matrix $\mathbf{A}_{Q}$), resulting in six potential attack scenarios, categorized into two types.
In a Type I attack, the adversary can query the target model with query graphs that come from the same distribution as the target’s training data. It is applicable in scenarios like side-effects prediction in drug interactions, financial fraud detection, and recommendation systems due to the widespread availability of such graphs.
In a Type II attack, the adversary attempts to steal the target model while lacking graph structural information. Therefore, the adversary must first try to learn the missing graph structure. Thus, in social networks, where the relationships between private user accounts are undisclosed, adversaries can only gather publicly available data from users’ profiles, lacking any insight into their social connections  [38].
@@ 4.2 Learning the Missing Graph Structure
In Type I attacks, the adversary can train the surrogate model $\mathcal{M}_{S}$ using responses from the target model to the provided query graph $\mathbf{G}_{Q}$. Conversely, Type II attacks necessitate the adversary to initially acquire knowledge of the absent graph structure $\mathbf{A}_{Q}$. Following the methodology outlined in [30], we employ the IDGL framework introduced by Chen et al. [3] for learning the query graph $\mathbf{G}_{Q}$. This involves minimizing a joint loss function that integrates both task-dependent prediction loss and graph regularization loss. To initiate the process, the adversary constructs a k-nearest neighbors algorithm ($k$NN) graph based on multi-head weighted cosine similarity. Subsequently, utilizing the IDGL framework, the adversary employs a joint loss function to explore and enhance the initial $k$NN graph structure by uncovering a hidden graph structure. This exploration is conducted by minimizing the joint loss function as detailed in [3]. Importantly, this reconstruction process is executed solely at the attacker’s end, without any interaction with the target model.
@@ 4.3 Training the Surrogate Model
With knowledge of the graph structure $\mathbf{A}_{Q}$, the adversary’s objective is to train a surrogate model $\mathcal{M}_{s}$. This model utilizes all nodes’ $l$-hop subgraphs from $\mathbf{G}_{Q}$ as training data and the query response $\mathbf{R}$ as supervised information. [30] observe that nodes that are close or connected in the query graph $\mathbf{G}_{Q}$ should be close in $\boldsymbol{\Upsilon}$ or $\mathbf{H}$, and the nodes that are of the same label should be close in $\boldsymbol{\Theta}$. Thus, it is possible to ignore the response types and uniformly treat all three possible target model responses as embedding vectors. The goal of training the surrogate model is to minimize the loss ($\mathcal{L}_{R}$) between $\hat{\mathbf{H}}_{Q}$ and $\mathbf{R}$. Notably, in this approach, the surrogate model’s output cannot be directly used for node classification tasks. After the surrogate model is trained and frozen an added MLP classification head ($\mathcal{O}$) is optimized, taking the surrogate model’s output ($\hat{\mathbf{H}}_{Q}$) as input and $\mathbf{C}_{Q}$ as supervision to minimize prediction error to facilitate node classification.
@@ 5.2 Spectral Graph Augmentations
In more detail, we consider the graph structure of the GNN training data to develop an optimal augmentation strategy. During surrogate model training, we incorporate spectral graph augmentations initially introduced by [19]. The augmented graph is constructed by extracting information with varying frequencies from the original graph.
@@
Such a pair of augmentations is termed an optimal contrastive pair.
Following this rule, we transform the adjacency matrix $A$ to a new augmentation $A^{\prime}$, where $A$ and $A^{\prime}$ form an optimal contrastive pair $Optim_{A,A^{\prime}}$ for all $\lambda_{m}\in[1,2]$ and $\lambda_{n}\in[0,1]$. The new graph, with its adjacency matrix $A^{\prime}$, undergoes further augmentations such as feature masking, edge dropping, and edge perturbation. The result serves as the input for training the surrogate model.
This form of augmentation improves the learning process of the surrogate model by boosting the effects of contrastive graph learning.

### v4

@@ Abstract
Graph Neural Networks (GNNs) are recognized as potent tools for processing real-world data organized in graph structures. Especially inductive GNNs, which allow for the processing of graph-structured data without relying on predefined graph structures, are becoming increasingly important in a wide range of applications. As such these networks become attractive targets for model-stealing attacks where an adversary seeks to replicate the functionality of the targeted network. Significant efforts have been devoted to developing model-stealing attacks that extract models trained on images and texts. However, little attention has been given to stealing GNNs trained on graph data. This paper identifies a new method of performing unsupervised model-stealing attacks against inductive GNNs, utilizing graph contrastive learning and spectral graph augmentations to efficiently extract information from the targeted model. The new type of attack is thoroughly evaluated on six datasets and the results show that our approach outperforms the current state-of-the-art by Shen et al. (2021). In particular, our attack surpasses the baseline across all benchmarks, attaining superior fidelity and downstream accuracy of the stolen model while necessitating fewer queries directed toward the target model. [^1] Code available at https://github.com/m-podhajski/EfficientGNNStealing
@@ 1 Introduction
Graph Neural Networks (GNNs) are specifically designed to operate on graph-structured data, such as molecules, social networks or complex infrastructure [8, 41, 28]. Unlike models developed for images or text, GNNs simultaneously consider both node features and graph structures. GNNs have garnered significant attention for their effectiveness in tasks involving structured data. A major advantage of GNNs is their ability to scale to large graphs while preserving computational efficiency. However, it is also well-documented that Machine Learning (ML) models, including GNNs, are vulnerable to model stealing attacks [34, 14, 6], where an adversary with query access to a target (victim) model can extract its parameters or functionality, often at a fraction of the cost required to train the model from scratch.
In such scenarios, the adversary sends a series of queries as inputs to the target model’s API and obtains the corresponding outputs. A local surrogate model is then trained to mimic the outputs of the target model for the same input data. Consequently, the surrogate model not only compromises the intellectual property of the target model but also potentially enables further attacks, such as extracting private training data [22] or constructing adversarial examples [26].
@@
Figure 1: Existing model stealing methods rely on aligning only the predictions for individual graph nodes sent to the victim and surrogate model. The new method identified in this paper greatly increases the amount of information extracted from the victim model outputs. We align the node representation from the surrogate model with the corresponding representation from the victim model while simultaneously distinguishing it from the representation of other nodes. By integrating spectral augmentations to augment the graph inputs of the surrogate models during its training process, we further increase the efficiency of the stealing process per query sent to the victim model.
@@
It is worth stressing that the majority of the current efforts on model stealing attacks concentrate on ML models for images and text data [34, 24, 37, 18]. Despite the growing significance of GNNs, model-stealing attacks against these types of neural networks are severely underexplored compared to image or text models. To our knowledge, only one study [33] addresses attacks on inductive GNNs, i.e. a type of model that can generalize well to unseen nodes. The authors define the threat model for model stealing attacks on inductive GNNs and describe six attack methods based on the adversary’s background knowledge and the responses of the target models. During the attack, the response of the surrogate model is aligned with the response of the target model by minimizing the RMSE loss between them. Although their approach yields promising results, the underlying method of training the surrogate model does not efficiently utilize the information available from the target’s responses. This increases the number of queries required to extract the target model’s API which results in higher chances of attack detection and mitigation.
@@
Inspired by recent approaches to encoder model stealing [6, 21, 31] we leverage the significant fact that augmented graph inputs should produce similar model outputs. Thus, the GNN model-stealing attack identified in this paper utilizes a contrastive objective at the node level. Specifically, for an input graph, we generate two graph views—one from the target model and one from the surrogate model for an augmented graph input. The surrogate model is trained by maximizing the agreement of node representations in these two views. This contrastive objective considers both positive and negative pairs between the node representations from the target and surrogate model. To augment the graph input to the surrogate model we leverage graph transformation operations derived via graph spectral analysis, enabling contrastive learning to capture useful structural representations. Those transformations include spectral graph cropping and graph frequency components reordering. Thus, the proposed methods that combine contrastive learning loss and effective graph-specific augmentations enable efficient usage of the information from the target model outputs. As a result, our approach fits into the framework established by Shen et al. (2021) [33]. It presents an enhanced surrogate model training methodology that significantly augments the efficiency and performance of the model-stealing process, outperforming the state-of-the-art method with less than half of its query budget.
@@
We evaluate all proposed attacks on three popular inductive GNN models including GraphSAGE [11], Graph Attention Network (GAT) [36], and Graph Isomorphism Network (GIN) [40] with six benchmark datasets, i.e., DBLP [25], Pubmed [30], Citeseer Full [9], Coauthor Physics [32], ACM [38], and Amazon Co-purchase Network for Photos [23]. The experiments demonstrate that our model-stealing attack consistently outperforms the state-of-the-art baseline approach across the vast majority of datasets and models. In the particular case of using SAGE as the surrogate model the increase in accuracy reached 36.8 %pt. for a GIN target trained on Amazon Co-purchase Network for Photos.
@@
  • We identify a new unsupervised GNN model stealing attack trained with a contrastive objective at the node level that compels the embedding of corresponding nodes from the surrogate and victim models to align while also distinguishing them from embeddings of other nodes within the surrogate and the target model.
  • We adapt spectral graph augmentations to generate multiple graph views for the surrogate model in the contrastive learning process.
  • We perform an extensive comparison of the newly identified stealing technique to the state of the art. The new stealing process demonstrates better performance while requiring fewer queries sent to the target model.
@@ 2 Preliminaries
$\mathbf{A^{\prime}}\in\langle-1,1\rangle^{n\times n}$ augmented adjacency matrix
@@
$u\in\mathcal{N}(v)$. In effect, the node embedding at the $k$-th layer captures the information within its $k$-hop neighborhood.
@@ 3 Threat Model
This section outlines the threat model, detailing the attack setting, adversary’s goals, and capabilities. The proposed approach is based on the threat model widely presented in [33].
Attack Setting. We investigate a challenging black-box scenario where the adversary is entirely unaware of the target GNN model’s specifics, including its parameters and architecture, and cannot interfere with the training process or the training graph $\mathbf{G}_{O}$. Our study focuses on the responses to three distinct types of node-level queries. In this context, the target model is an inductive GNN that utilizes the $l$-hop subgraph $\mathbf{G}_{v}^{l}$ of a node $v$ as input to generate outputs for that node. These outputs could be in the form of predicted posterior probabilities, node embedding vectors, or a 2-dimensional t-SNE projection.
Adversary’s Goal. According to the classification established by [14], adversaries typically have two primary objectives: theft or reconnaissance. A theft adversary aims to create a surrogate model $\mathcal{M}_{S}$ that replicates the performance of the target model $\mathcal{M}_{T}$ for the specific task at hand [34, 26]. On the other hand, a reconnaissance adversary seeks to develop a surrogate model $\mathcal{M}_{S}$ that approximates $\mathcal{M}_{T}$ across all possible inputs. Achieving such a high level of fidelity allows the adversary to use the surrogate model for subsequent attacks, such as generating adversarial examples, without needing to interact directly with the target model $\mathcal{M}_{T}$ [26].
Adversary’s Capabilities. Firstly, we consider a scenario where the adversary interacts with a target model $\mathcal{M}_{T}$ via a public API [34, 24, 12, 13]. The adversary submits a query graph $\mathbf{G}_{Q}$ and receives responses $\mathbf{R}$, which could be a node embedding matrix ($\mathbf{H}$), a predicted probability matrix ($\boldsymbol{\Theta}$), or a t-SNE projection matrix ($\boldsymbol{\Upsilon}$). These responses represent typical outputs from such APIs, varying in the level of information they provide for tasks like graph visualization, transfer learning, and fine-tuning GNNs.
Additionally, We assume the query graph $\mathbf{G}_{Q}$, with node features $\mathbf{X}_{Q}$ and graph structure $\mathbf{A}_{Q}$, is drawn from the same distribution as the training graph $\mathbf{G}_{O}$ used to train $\mathcal{M}_{T}$. This implies that $\mathbf{G}_{Q}$ and $\mathbf{G}_{O}$ are sampled from the same dataset, consistent with studies on neural network attacks using public datasets [14, 12], particularly in graph-intensive domains like social networks and molecular structures. As discussed in Section 4.2, this assumption can be adjusted.
@@ 4 Attack Framework
This section aims to discuss attack scenarios that can be launched by the adversary depending on different levels of knowledge and describe the general framework of GNN model-stealing. We adopt both, the attack structure and taxonomy proposed by Shen and co-authors [33].
@@ 4.1 Attack Taxonomy
To conduct a model stealing attack, the adversary relies on two essential components: the query graph $\mathbf{G}_{Q}=(\mathbf{A}_{Q},\mathbf{X}_{Q},\mathbf{C}_{Q})$ and the corresponding response $\mathbf{R}$. The query graph consists of
  • $\mathbf{A}_{Q}$: the adjacency matrix, which represents the graph structure,
  • $\mathbf{X}_{Q}$: the node feature matrix, detailing the attributes of each node,
  • $\mathbf{C}_{Q}$: the labels or class information associated with nodes or edges.
The response $\mathbf{R}$ from the target model can be in one of several forms, such as
  • $\mathbf{H}$: node embeddings, which are vectors representing each node in the graph,
  • $\boldsymbol{\Theta}$: predicted posterior probabilities, indicating the model’s confidence in various outcomes,
  • $\boldsymbol{\Upsilon}$: a t-SNE projection matrix, which visualizes node embeddings in a 2-dimensional space.
The adversary may not have complete structural information about the query graph $\mathbf{G}_{Q}$, lacking the adjacency matrix $\mathbf{A}_{Q}$, which leads to six possible attack scenarios grouped into two types. In a Type I attack, the adversary queries the target model using graphs from the same distribution as those used in training. This is common in domains like drug interaction predictions, fraud detection, and recommendation systems, where abundant graph data is available.
Conversely, a Type II attack involves compromising the model without access to the specific graph structure used during training. Here, the adversary must reconstruct or infer missing information, such as in social networks where user connections are private, and only public profile information is accessible [42].
@@ 4.2 Learning the Missing Graph Structure
In Type I attacks, the adversary trains a surrogate model $\mathcal{M}_{S}$ using responses from the target model to the query graph $\mathbf{G}_{Q}$. In contrast, Type II attacks require inferring the missing graph structure $\mathbf{A}_{Q}$. Following [33], we use the IDGL framework by Chen et al. [3] to infer $\mathbf{G}_{Q}$. This involves minimizing a joint loss function that combines task-specific prediction loss with graph regularization. The adversary starts with a k-nearest neighbors ($k$NN) graph based on multi-head weighted cosine similarity and iteratively improves it using IDGL, optimizing the joint loss function without interacting directly with the target model.
@@ 4.3 Training the Surrogate Model
When the adversary has access to the graph structure $\mathbf{A}_{Q}$, their goal is to train a surrogate model $\mathcal{M}_{S}$. This model is trained using the $l$-hop subgraphs of all nodes from the query graph $\mathbf{G}_{Q}$ as input and the corresponding responses $\mathbf{R}$ from the target model as supervised signals. As noted by [33], nodes that are close or connected in $\mathbf{G}_{Q}$ should also be close in the t-SNE projection matrix $\boldsymbol{\Upsilon}$ or the node embedding matrix $\mathbf{H}$. Thus, the target model’s responses can uniformly be treated as embedding vectors. The primary goal is to minimize the loss ($\mathcal{L}_{R}$) between the surrogate model’s output embeddings $\hat{\mathbf{H}}_{Q}$ and the target model’s responses $\mathbf{R}$. It’s important to note that the surrogate model’s output cannot be directly used for node classification, as it is trained to mimic the target model’s responses, not to perform classification. An additional step is required to enable node classification. After training and freezing the surrogate model, an MLP classification head ($\mathcal{O}$) is optimized, using the surrogate model’s output ($\hat{\mathbf{H}}_{Q}$) as input and $\mathbf{C}_{Q}$ as supervision to minimize prediction error. This two-step process ensures the surrogate model’s outputs are effectively utilized for node classification.
@@ 5.2 Spectral Graph Augmentations
Specifically, we consider the graph structure of the GNN training data to develop an optimal augmentation strategy. During surrogate model training, we incorporate spectral graph augmentations initially introduced by [20]. The augmented graph is constructed by extracting information with varying frequencies from the original graph.
@@
Such a pair of augmentations is termed an optimal contrastive pair as in [20]. On a high level, ensuring that the difference in high-frequency components is greater than in low-frequency components helps the model to learn more robust and discriminative features.
Following this rule, we transform the adjacency matrix $A$ to a new augmentation $A^{\prime}$, where $A$ and $A^{\prime}$ form an optimal contrastive pair $Optim_{A,A^{\prime}}$ for all $\lambda_{m}\in[1,2]$ and $\lambda_{n}\in[0,1]$. The new graph, with its adjacency matrix $A^{\prime}$, undergoes further augmentations such as feature masking, edge dropping, and edge perturbation. The result serves as the input for training the surrogate model. This form of augmentation improves the learning process of the surrogate model by boosting the effects of contrastive graph learning.